OpenStudy (anonymous):

Multiply (2 – 7i)(9 + 5i) somebody please help me :)

OpenStudy (anonymous):

53-53i

OpenStudy (anonymous):

Treat it as variable first,make it easier sort of

OpenStudy (dumbcow):

= (2*9) + (2*5i) + (-7i*9) + (-7i*5i) = 18+10i-63i-35i^2 ---> i^2 = -1 = 18-53i -35(-1) =18 -53i +35 =53 -53i